A particle is moving with the given data. Find the position of the particle.

a(t) = t’ – 4t + 6, s(0) = 0, s(1) = 20
Answer #1

aC=+? 47+6 :: ACH=CVCH) Sacvct)=facher VCt={[4241+6 ]cut VCH= x - 21+6H+C1 • VCH)=d(SCH)) de Sacsch)=fvcudo SCHS 28+66+6) SCH